# A Comprehensive Survey on Knowledge Distillation

### v1

@@ VII Performance Comparison
FitNet [22] (2014) Feature WRN-40-2 (75.61) WRN-40-1 (71.98) 72.24 (+0.26) Link
KD [10] (2015) Logit WRN-40-2 (75.61) WRN-40-1 (71.98) 73.54 (+1.56) Link
AT [23] (2016) Feature WRN-40-2 (75.61) WRN-40-1 (71.98) 72.77 (+1.79) Link
CCKD [86] (2019) Similarity ResNet-110 (–) ResNet-20 (68.40) 72.40 (+4.00)
SPKD [25] (2019) Similarity WRN-40-2 (75.61) WRN-40-1 (71.98) 72.43 (+1.45)
Chen et al. [48] (2020) Similarity + Logit ResNet-101 (71.77) ResNet-18 (65.64) 69.14 (+3.49) Link
TAKD [18] (2020) Logit ResNet-110 (–) ResNet-8 (–) 61.82 Link
SphericalKD [29] (2020) Logit ResNet-32$\times$4 (79.42) ResNet-8$\times$4 (72.50) 76.40 (+3.90) Link
RKD [68] (2021) Feature WRN-40-2 (75.61) WRN-40-1 (71.98) 75.09 (+3.11) Link
SemCKD [69] (2021) Feature + Logit WRN-40-2 (75.61) MobileNetV2 (65.43) 69.67 (+4.24) Link
ICKD [24] (2021) Similarity + Logit WRN-40-2 (75.61) WRN-40-1 (71.98) 74.63 (+2.65) Link
DKD [35] (2022) Logit WRN-40-2 (75.61) WRN-40-1 (71.98) 74.81 (+2.83) Link
DistKD [55] (2022) Similarity + Logit WRN-40-2 (75.61) WRN-40-1 (71.98) 74.73 (+2.75) Link
Norm [79] (2023) Feature WRN-40-2 (75.61) WRN-40-1 (71.98) 74.82 (+2.84) Link
SFKD [34] (2024) Logit ResNet-32$\times$4 (79.55) ResNet-8$\times$4 (72.50) 76.84 (+4.34)
NormKD [30] (2024) Logit ResNet-32$\times$4 (79.42) ResNet-8$\times$4 (72.50) 76.57 (+4.07) Link
CRLD [63] (2024) Similarity + Logit WRN-40-2 (75.61) WRN-40-1 (71.98) 75.58 (+3.60) Link
NTCE-KD [36] (2024) Logit WRN-40-2 (75.61) WRN-40-1 (71.98) 76.44 (+4.46)
TTM [32] (2024) Logit WRN-40-2 (75.61) WRN-40-1 (71.98) 74.58 (+2.60) Link
Miles et al. [31] (2024) Logit WRN-40-2 (76.46) WRN-16-2 (73.64) 76.14 (+2.50) Link
SDD [39] (2024) Logit ResNet-32$\times$4 (79.42) MobileNetV2 (64.6) 68.84 (+4.24) Link
Sun et al. [20] (2024) Logit WRN-40-2 (75.61) ResNet-8$\times$4 (72.50) 77.11 (+3.14) Link
Segmentation (CityScapes)
Method Knowledge Teacher (baseline) Student (baseline) mIoU Code
KD [10] (2015) Logit ResNet-101 (77.66) ResNet-18 (64.09) 65.21 (+1.12) Link
AT [23] (2016) Feature ResNet-101 (77.66) ResNet-18 (64.09) 65.29 (+1.20) Link
Xie et al. [84] (2018) Similarity + Logit ResNet-101 (70.90) MobileNetV2 (67.30) 71.90 (+4.60)
SKD [58] (2019) Feature + Logit ResNet-101 (78.40) MobileNetV2 (67.60) 71.40 (+3.80) Link
He et al. [82] (2019) Feature + Similarity ResNet-101 (71.40) MobileNetV2 (70.20) 72.70 (+2.50)
IFVD [26] (2020) Similarity + Logit ResNet-101 (78.56) ResNet-18 (69.10) 74.54 (+5.44) Link
CWD [67] (2021) Feature ResNet-101 (78.50) ResNet-18 (63.63) 71.03 (+7.40) Link
DSD [54] (2021) Similarity ResNet-101 (78.23) ResNet-18 (69.42) 72.24 (+2.82)
CIRKD [59] (2022) Similarity + Logit ResNet-101 (78.07) ResNet-18 (72.55) 74.73 (+2.18) Link
DIST [55] (2022) Similarity + Logit ResNet-101 (78.07) ResNet-18 (72.55) 76.31 (+3.76) Link
IDD [56] (2022) Similarity ResNet-101 (78.40) MobileNetV2 (67.60) 76.33 (+8.73)
MGD [70] (2022) Feature ResNet-101 (78.34) ResNet-18 (73.20) 76.31 (+3.11) Link
MLP [71] (2023) Feature ResNet-101 (78.34) ResNet-18 (73.20) 76.55 (+3.35)
DiffKD [72] (2023) Feature ResNet-101 (78.07) ResNet-18 (74.21) 77.78 (+3.57) Link
FAKD [77] (2024) Feature ResNet-101 (79.74) ResNet-18 (68.99) 74.75 (+5.76)
BPKD [40] (2024) Logit ResNet-101 (79.74) ResNet-18 (74.23) 77.57 (+3.34) Link
LAD [76] (2024) Feature + Logit ResNet-101 (79.76) ResNet-18 (72.65) 76.86 (+4.21)
FreeKD [78] (2024) Feature ResNet-101 (70.43) ResNet-18 (73.20) 76.45 (+3.25) Link
AttnFD [74] (2024) Feature ResNet-101 (77.66) ResNet-18 (64.09) 73.04 (+8.96) Link
AICSD [57] (2025) Similarity + Logit ResNet-101 (77.66) ResNet-18 (64.09) 70.96 (+6.88) Link
@@
TABLE XIII: Performance comparison of different Black-box and White-box distillation methods for LLMs (* denotes that the comparison is between the student model before and after distillation).
@@
DistillBERT [453] (2019) Logit $BERT_{base}$ 1.66 GLUE 77.00 / 79.50 97.0%
TinyBERT [452] (2019) Feature $BERT_{base}$ 7.50 GLUE 77.00 / 79.50 97.0% Link
PKD [454] (2019) Logit + Feature $BERT_{base}$ 2.00 GLUE 77.70 / 84.90 92.0% Link
PD [455] (2019) Logit $BERT_{base}$ 2.00 GLUE 82.10 / 81.70 100.5% Link
Xtremedistil [458] (2020) Feature $mBERT_{base}$ 35.00 Multilingual NER 88.60 / 92.70 95.0% Link
MobileBERT [456] (2020) Feature $BERT_{base}$ 4.30 GLUE 77.70 / 78.30 99.0% Link
MINILM [457] (2020) Feature $BERT_{base}$ 1.65 GLUE 80.40 / 81.50 98.0%
xtremedistiltransformers [459] (2021) Feature $BERT_{base}$ 7.70 GLUE 81.70 / 83.70 97.6% Link
MetaDistil [460] (2021) Feature $BERT_{base}$ 2.00 GLUE 80.40 / 80.70 99.0% Link
TED [461] (2023) Feature $DeBERTaV3_{base}$ 2.60 GLUE 87.50 / 88.90 98.0% Link
MiniLLM [465] (2024) Feature GPT-2 2.00 DollyEval 57.40 / 58.40 94.0% Link
PC-LoRA [464] (2024) Feature $BERT_{base}$ 3.73 IMDB 92.78 / 94.00 98.0%
GKD [463] (2024) Logit $T5_{XL}$ 3.75 GSM8K 29.10 / 29.00 100.3%

### v2

@@ 7 Performance Comparison
FitNet (Romero et al., 2014) Feature WRN-40-2 (75.61) WRN-40-1 (71.98) 72.24 (+0.26) Link
[…]
KD (Hinton, 2015) Logit WRN-40-2 (75.61) WRN-40-1 (71.98) 73.54 (+1.56) Link
[…]
AT (Zagoruyko & Komodakis, 2016) Feature WRN-40-2 (75.61) WRN-40-1 (71.98) 72.77 (+1.79) Link
[…]
CCKD (Peng et al., 2019) Similarity ResNet-110 (–) ResNet-20 (68.40) 72.40 (+4.00)
SPKD (Tung & Mori, 2019) Similarity WRN-40-2 (75.61) WRN-40-1 (71.98) 72.43 (+1.45)
[…]
CRD (Tian et al., 2019) Feature WRN-40-2 (75.61) WRN-40-1 (71.98) 74.14 (+2.16) Link
[…]
OFD (Heo et al., 2019a) Feature WRN-40-2 (75.61) WRN-40-1 (71.98) 74.33 (+2.35) Link
[…]
Chen et al. (Chen et al., 2020f) Similarity + Logit ResNet-101 (71.77) ResNet-18 (65.64) 69.14 (+3.49) Link
TAKD (Mirzadeh et al., 2020) Logit ResNet-110 (–) ResNet-8 (–) 61.82 Link
[…]
RKD (Chen et al., 2021d) Feature WRN-40-2 (75.61) WRN-40-1 (71.98) 75.09 (+3.11) Link
[…]
SemCKD (Chen et al., 2021a) Feature + Logit WRN-40-2 (75.61) MobileNetV2 (65.43) 69.67 (+4.24) Link
[…]
ICKD (Liu et al., 2021c) Similarity + Logit WRN-40-2 (75.61) WRN-40-1 (71.98) 74.63 (+2.65) Link
DKD (Zhao et al., 2022a) Logit WRN-40-2 (75.61) WRN-40-1 (71.98) 74.81 (+2.83) Link
[…]
DistKD (Huang et al., 2022a) Similarity + Logit WRN-40-2 (75.61) WRN-40-1 (71.98) 74.73 (+2.75) Link
[…]
Norm (Liu et al., 2023a) Feature WRN-40-2 (75.61) WRN-40-1 (71.98) 74.82 (+2.84) Link
[…]
CRLD (Zhang et al., 2024g) Similarity + Logit WRN-40-2 (75.61) WRN-40-1 (71.98) 75.58 (+3.60) Link
[…]
NTCE-KD (Li et al., 2024a) Logit WRN-40-2 (75.61) WRN-40-1 (71.98) 76.44 (+4.46)
[…]
TTM (Zheng & Yang, 2024) Logit WRN-40-2 (75.61) WRN-40-1 (71.98) 74.58 (+2.60) Link
[…]
Miles et al. (Miles & Mikolajczyk, 2024) Logit WRN-40-2 (76.46) WRN-16-2 (73.64) 76.14 (+2.50) Link
[…]
ResNet50 (79.34) MobileNetV2 (64.60) 71.46 (+6.86)
[…]
InDistill (Sarridis et al., 2025) Feature WRN-40-2 (75.61) WRN-40-1 (71.98) 75.09 (+3.11) Link
IKD (Wang et al., 2025a) Logit WRN-40-2 (75.61) WRN-40-1 (71.98) 74.84 (+2.86)
ResNet50 (79.34) MobileNetV2 (64.60) 70.49 (+5.89)
ADG-KD (Li et al., 2025b) Logit WRN-40-2 (75.61) WRN-40-1 (71.98) 75.84 (+3.86)
[…]
TeKAP (Hossain et al., 2025) Feature + Logit WRN-40-2 (75.61) WRN-40-1 (71.98) 74.41 (+2.43) Link
[…]
In this work, existing methods are compared for two well-known tasks: classification and semantic segmentation segmentation, as almost all the proposed distillation methods evaluate their results on these two tasks. For classification, the ImageNet and CIFAR-100 datasets, and for semantic segmentation, the Cityscapes and PascalVOC datasets are used, as these two datasets are widely employed in existing methods.
ImageNet is a large-scale dataset designed for visual object recognition tasks, consisting of over 1.2 million training images, 50,000 validation images, and 100,000 testing images across 1,000 object classes. CIFAR-100 is composed of $32\times 32$ images taken from 100 classes, with 50,000/10,000 images for training/testing, and each class has the same number of training and testing images. Cityscapes is designed for understanding urban scenes and includes 2,975/500/1,525 images for training/validation/testing in 19 classes. PascalVOC dataset comprises 1464/1449/1456 images for train/val/test, with 21 classes.
Table 13: Performance comparison of different knowledge distillation methods on the ImageNet classification dataset (* denotes that results are not reported from the original paper).
FitNet (Romero et al., 2014) Logit ResNet-34 (73.31) ResNet-18 (69.75) 70.31 (+0.56) Link
KD (Hinton, 2015) Logit ResNet-34 (73.31) ResNet-18 (69.75) 70.62 (+0.87) Link
ResNet-50 (76.16) MobileNetV1 (68.87) 68.58 (-0.29)
AT (Zagoruyko & Komodakis, 2016) Feature ResNet-34 (73.31) ResNet-18 (69.75) 70.30 (+0.55) Link
ResNet-50 (76.16) MobileNetV1 (68.87) 69.56 (+0.69)
CRD (Tian et al., 2019) Feature ResNet-34 (73.31) ResNet-18 (69.75) 71.17 (+1.42) Link
ResNet-50 (76.16) MobileNetV1 (68.87) 71.37 (+2.50)
OFD (Heo et al., 2019a) Feature ResNet-34 (73.31) ResNet-18 (69.75) 70.81 (+1.06) Link
ResNet-50 (76.16) MobileNetV1 (68.87) 71.25 (+2.38)
SphericalKD (Guo et al., 2020a) Logit ResNet-34 (73.31) ResNet-18 (69.75) 72.30 (+2.55) Link
RKD (Chen et al., 2021d) Feature + Logit ResNet-34 (73.31) ResNet-18 (69.75) 71.61 (+1.86) Link
ResNet-50 (76.16) MobileNetV1 (68.87) 72.56 (+3.69)
SemCKD (Chen et al., 2021a) Feature + Logit ResNet-34 (73.31) ResNet-18 (69.75) 70.87 (+1.12) Link
ICKD (Liu et al., 2021c) Similarity + Logit ResNet-34 (73.31) ResNet-18 (69.75) 72.19 (+2.44) Link
DKD (Zhao et al., 2022a) Logit ResNet-34 (73.31) ResNet-18 (69.75) 72.05 (+2.30) Link
ResNet-50 (76.16) MobileNetV1 (68.87) 72.05 (+3.18)
DistKD (Huang et al., 2022a) Similarity + Logit ResNet-34 (73.31) ResNet-18 (69.75) 72.07 (+2.32) Link
ResNet-50 (76.16) MobileNetV1 (70.13) 73.24 (+3.11)
CAT-KD (Guo et al., 2023d) Feature ResNet-34 (73.31) ResNet-18 (69.75) 71.26 (+1.51) Link
ResNet-50 (76.16) MobileNetV1 (70.13) 72.24 (+3.37)
NKD (Yang et al., 2023g) Logit ResNet-34 (73.31) ResNet-18 (69.75) 71.96 (+2.21) Link
ResNet-50 (76.16) MobileNetV1 (70.13) 72.58 (+3.71)
MLLD (Jin et al., 2023a) Logit ResNet-34 (73.31) ResNet-18 (69.75) 71.90 (+2.15) Link
ResNet-50 (76.16) MobileNetV1 (70.13) 73.01 (+4.14)
Norm (Liu et al., 2023a) Feature ResNet-34 (73.31) ResNet-18 (69.75) 72.14 (+2.39) Link
ResNet-50 (76.16) MobileNetV1 (68.87) 74.26 (+5.39)
SFKD (Yuan et al., 2024b) Logit ResNet-34 (73.31) ResNet-18 (69.75) 71.86 (+2.11)
ResNet-50 (76.16) MobileNetV1 (68.87) 72.24 (+3.37)
NormKD (Chi et al., 2023) Logit ResNet-34 (73.31) ResNet-18 (69.75) 72.03 (+2.28) Link
ResNet-50 (76.16) MobileNetV1 (68.87) 72.79 (+3.92)
CRLD (Zhang et al., 2024g) Similarity + Logit ResNet-34 (73.31) ResNet-18 (69.75) 72.37 (+2.62) Link
ResNet-50 (76.16) MobileNetV1 (70.13) 73.53 (+4.66)
NTCE-KD (Li et al., 2024a) Logit ResNet-34 (73.31) ResNet-18 (69.75) 73.12 (+3.37)
ResNet-50 (76.16) MobileNetV1 (70.13) 74.90 (+6.03)
TTM (Zheng & Yang, 2024) Logit ResNet-34 (73.31) ResNet-18 (69.75) 72.19 (+2.44) Link
ResNet-50 (76.16) MobileNetV1 (70.13) 773.09 (+4.22)
SDD (Wei et al., 2024b) Logit ResNet-34 (73.31) ResNet-18 (69.75) 72.33 (+2.58) Link
Sun et al. (2024a) Logit ResNet-34 (73.31) ResNet-18 (69.75) 72.08 (+2.33) Link
ResNet-50 (76.16) MobileNetV1 (68.87) 73.22 (+4.35)
InDistill (Sarridis et al., 2025) Feature ResNet-34 (73.31) ResNet-18 (69.75) 71.34 (+1.59) Link
IKD (Wang et al., 2025a) Logit ResNet-34 (73.31) ResNet-18 (69.75) 71.82 (+2.07)
ADG-KD (Li et al., 2025b) Logit ResNet-34 (73.31) ResNet-18 (69.75) 72.22 (+2.47)
ResNet-50 (76.16) MobileNetV1 (68.87) 73.43 (+4.56)
Tables 12 and 13 summarize various knowledge distillation methods applied to classification tasks on the CIFAR-100 and ImageNet datasets. For each method, the source of knowledge, teacher and student architectures, and post-distillation accuracy are reported. In most cases, two scenarios are considered: one where the teacher and student share similar architectures, and another where their architectures differ.
Table 14: Performance comparison of different knowledge distillation methods on the PascalVoc segmentation dataset (* denotes that results are not reported from the original paper).
FitNet (Romero et al., 2014) Logit DeepLab-R50 (76.21) DeepLab-MV2 (70.57) 71.30 (+0.73) Link
KD (Hinton, 2015) Logit DeepLab-R101 (77.85) DeepLab-R18 (67.50) 69.13 (+1.63) Link
DeepLab-R101 (77.85) DeepLab-MV2 (63.92) 66.39 (+2.47)
AT (Zagoruyko & Komodakis, 2016) Feature DeepLab-R101 (77.85) DeepLab-R18 (67.50) 68.95 (+1.09) Link
DeepLab-R101 (77.85) DeepLab-MV2 (63.92) 66.27 (+2.35)
SKD (Liu et al., 2019d) Feature + Logit PSPNet-R101 (78.52) PSPNet-R18 (71.35) 72.01 (+0.66) Link
PSPNet-R101 (77.67) DeepLab-R18 (71.98) 73.03 (+1.05)
He et al. (2019) Feature + Similarity DeepLab-R50 (76.21) DeepLab-MV2 (70.57) 72.50 (+1.93)
IFVD (Wang et al., 2020d) Similarity + Logit PSPNet-R101 (78.52) PSPNet-R18 (71.35) 73.49 (+2.14) Link
PSPNet-R101 (77.67) DeepLab-R18 (71.98) 72.87 (+0.89)
CWD (Shu et al., 2021) Feature PSPNet-R101 (78.52) PSPNet-R18 (71.35) 73.49 (+2.14) Link
PSPNet-R101 (77.67) DeepLab-R18 (71.98) 73.78 (+1.79)
DSD (Feng et al., 2021) Similarity DeepLab-R101 (79.10) DeepLab-R18 (71.9) 73.70 (+1.80)
CIRKD (Yang et al., 2022c) Similarity + Logit DeepLab-R101 (77.67) DeepLab-R18 (73.21) 74.50 (+1.29) Link
DeepLab-R101 (77.67) PSPNet-R18 (73.33) 74.78 (+1.45)
DIST (Huang et al., 2022a) Similarity + Logit DeepLab-R101 (77.85) DeepLab-R18 (67.50) 69.84 (+2.34) Link
DeepLab-R101 (77.85) PSPNet-R18(67.40) 68.93 (+1.53)
IDD (Zhang et al., 2022d) Similarity PSPNet-R101 (77.86) PSPNet-R18 (70.80) 76.70 (+5.9)
MGD (Yang et al., 2022f) Feature PSPNet-R101 (78.52) PSPNet-R18 (71.35) 74.98 (+3.63) Link
FAKD (Yuan et al., 2024a) Feature PSPNet-R101 (78.52) PSPNet-R18 (70.52) 73.97 (+3.45)
DeepLab-R101 (78.62) DeepLab-MV2 (62.56) 67.62 (+5.06)
BPKD (Liu et al., 2024d) Logit PSPNet-R101 (78.52) PSPNet-R18 (71.35) 75.74 (+4.39) Link
LAD (Liu et al., 2024g) Feature + Logit PSPNet-R101 (78.52) PSPNet-R18 (71.35) 75.74 (+4.39)
PSPNet-R101 (78.52) DeepLab-R18 (71.98) 76.33 (+4.35)
AttnFD (Mansourian et al., 2024) Feature DeepLab-R101 (77.85) DeepLab-R18 (67.50) 73.09 (+5.59) Link
DeepLab-R101 (77.85) PSPNet-R18(67.40) 70.95 (+3.55)
AICSD (Mansourian et al., 2025) Similarity + Logit DeepLab-R101 (77.85) DeepLab-R18 (67.50) 70.58 (+3.08) Link
DeepLab-R101 (77.85) PSPNet-R18(66.60) 68.29 (+1.69)
Table 15: Performance comparison of different knowledge distillation methods on the Cityscapes segmentation dataset (* denotes that results are not reported from the original paper).
KD (Hinton, 2015) Logit DeepLab-R101 (77.66) DeepLab-R18 (64.09) 65.21 (+1.12) Link
DeepLab-R101 (77.66) DeepLab-MV2 (63.05) 64.03 (+0.98)
AT (Zagoruyko & Komodakis, 2016) Feature DeepLab-R101 (77.66) DeepLab-R18 (64.09) 65.29 (+1.20) Link
DeepLab-R101 (77.66) DeepLab-MV2 (63.05) 63.72 (+0.67)
Xie et al. (2018) Similarity + Logit DeepLab-R101 (70.90) DeepLab-MV2 (67.30) 71.90 (+4.60)
SKD (Liu et al., 2019d) Feature + Logit PSPNet-R101 (79.76) PSPNet-R18 (72.65) 74.23 (+1.58) Link
PSPNet-R101 (79.76) DeepLab-R18 (74.96) 75.32 (+0.36)
He et al. (2019) Feature + Similarity DeepLab-R101 (71.40) DeepLab-MV2 (70.20) 72.70 (+2.50)
IFVD (Wang et al., 2020d) Similarity + Logit PSPNet-R101 (79.76) PSPNet-R18 (72.65) 74.55 (+1.90) Link
PSPNet-R101 (79.76) DeepLab-R18 (74.96) 76.01 (+1.05)
CWD (Shu et al., 2021) Feature PSPNet-R101 (79.76) PSPNet-R18 (72.65) 75.91 (+3.26) Link
PSPNet-R101 (79.76) DeepLab-R18 (74.96) 77.13 (+2.17)
DSD (Feng et al., 2021) Similarity DeepLab-R101 (78.23) DeepLab-R18 (69.42) 72.24 (+2.82)
CIRKD (Yang et al., 2022c) Similarity + Logit DeepLab-R101 (78.07) DeepLab-R18 (74.21) 76.38 (+2.27) Link
DeepLab-R101 (78.07) PSPNet-R18 (72.55) 74.73 (+2.18)
DIST (Huang et al., 2022a) Similarity + Logit DeepLab-R101 (78.07) DeepLab-R18 (74.21) 77.10 (+2.89) Link
DeepLab-R101 (78.07) PSPNet-R18 (72.55) 76.31 (+3.76)
IDD (Zhang et al., 2022d) Similarity PSPNet-R101 (78.50) PSPNet-R18 (70.09) 77.59 (+7.5)
PSPNet-R101 (78.50) ESPNet (61.40) 68.87 (+7.47)
MGD (Yang et al., 2022f) Feature PSPNet-R101 (78.34) PSPNet-R18 (69.85) 73.63 (+3.78) Link
PSPNet-R101 (78.34) DeepLab-R18 (73.20) 76.31 (+3.11)
MLP (Liu et al., 2023c) Feature PSPNe-R101 (78.34) PSPNet-R18 (69.85) 74.51 (+4.66)
PSPNet-R101 (78.34) DeepLab-R18 (73.20) 76.55 (+3.35)
DiffKD (Huang et al., 2023b) Feature DeepLab-R101 (78.07) DeepLab-R18 (74.21) 77.78 (+3.57) Link
DeepLab-R101 (78.07) PSPNet-R18 (72.55) 75.83 (+3.28)
FAKD (Yuan et al., 2024a) Feature PSPNet-R101 (79.74) PSPNet-R18 (68.99) 74.75 (+5.76)
DeepLab-R101 (80.98) DeepLab-MV2 (70.49) 74.08 (+3.59)
BPKD (Liu et al., 2024d) Logit PSPNet-R101 (79.74) PSPNet-R18 (74.23) 77.57 (+3.34) Link
DeepLab-R101 (80.98) DeepLab-MV2 (75.29) 78.59 (+3.30)
LAD (Liu et al., 2024g) Feature + Logit PSPNet-R101 (79.76) PSPNet-R18 (72.65) 76.86 (+4.21)
PSPNet-R101 (79.76) DeepLab-R18 (74.96) 77.23 (+2.27)
FreeKD (Zhang et al., 2024i) Feature PSPNet-R101 (78.34) PSPNet-R18 (69.85) 74.40 (+4.55) Link
PSPNet-R101 (78.34) DeepLab-R18 (73.20) 76.45 (+3.25)
AttnFD (Mansourian et al., 2024) Feature DeepLab-R101 (77.66) DeepLab-R18 (64.09) 73.04 (+8.96) Link
DeepLab-R101 (77.66) PSPNet-R18 (65.72) 68.86 (+3.14)
AICSD (Mansourian et al., 2025) Similarity + Logit DeepLab-R101 (77.66) DeepLab-R18 (64.09) 70.96 (+6.88) Link
DeepLab-R101 (77.66) DeepLab-MV2 (63.05) 69.51 (+6.56)
In a similar vein, Tables 14 and 15 report the results of various methods for the semantic segmentation task on the PASCAL VOC and Cityscapes datasets. For a fair comparison, the reported results are taken directly from the original papers. However, in cases where results on the specified datasets are not provided in the original work, we report results obtained from our own implementations (denoted by *).
Along the same line, Table 16 presents a comparison of distillation methods for LLMs. It shows the compression rate and the improvement over the teacher model for each method. As can be seen, distillation in LLMs is more crucial than in other fields, as it allows for compressing large models with a high compression rate while still maintaining comparable performance. It should be noted that in cases where the exact performance of the teacher model is not reported, performance of the student model before and after distillation is reported (denoted by *).
@@
Table 16: Performance comparison of different Black-box and White-box distillation methods for LLMs (* denotes that the comparison is between the student model before and after distillation).
@@
DistillBERT (Sanh, 2019) Logit $BERT_{base}$ 1.66 GLUE 77.00 / 79.50 97.0%
TinyBERT (Jiao et al., 2019) Feature $BERT_{base}$ 7.50 GLUE 77.00 / 79.50 97.0% Link
PKD (Sun et al., 2019) Logit + Feature $BERT_{base}$ 2.00 GLUE 77.70 / 84.90 92.0% Link
PD (Turc et al., 2019) Logit $BERT_{base}$ 2.00 GLUE 82.10 / 81.70 100.5% Link
Xtremedistil (Mukherjee & Awadallah, 2020) Feature $mBERT_{base}$ 35.00 Multilingual NER 88.60 / 92.70 95.0% Link
MobileBERT (Sun et al., 2020) Feature $BERT_{base}$ 4.30 GLUE 77.70 / 78.30 99.0% Link
MINILM (Wang et al., 2020b) Feature $BERT_{base}$ 1.65 GLUE 80.40 / 81.50 98.0%
xtremedistiltransformers (Mukherjee et al., 2021) Feature $BERT_{base}$ 7.70 GLUE 81.70 / 83.70 97.6% Link
MetaDistil (Zhou et al., 2021b) Feature $BERT_{base}$ 2.00 GLUE 80.40 / 80.70 99.0% Link
TED (Liang et al., 2023a) Feature $DeBERTaV3_{base}$ 2.60 GLUE 87.50 / 88.90 98.0% Link
MiniLLM (Gu et al., 2024) Feature GPT-2 2.00 DollyEval 57.40 / 58.40 94.0% Link
PC-LoRA (Hwang et al., 2024) Feature $BERT_{base}$ 3.73 IMDB 92.78 / 94.00 98.0%
GKD (Agarwal et al., 2024) Logit $T5_{XL}$ 3.75 GSM8K 29.10 / 29.00 100.3%